OpenStudy (anonymous):

Write the polynomial equation of lowest degree with real coefficients having 2 and (1-3i) as two of its roots.

hartnn (hartnn):

if a+bi is a root, then a-bi is also a root of that equation. (with real co-efficients) so,if here 1-3i is a root, whats the other one?

OpenStudy (anonymous):

@hartnn 1+3i.

hartnn (hartnn):

correct! now if a is the root of f(x), then (x-a) is one of the factor. so when '2' iss the root, the factor is ?

OpenStudy (anonymous):

@hartnn x-2.

hartnn (hartnn):

correct! now 1-3i is also a root, so (x-(1-3i)) = (x-1+3i) is also a factor, lastly, since 1+3i is also a root, which is the last factor of f(x) ?

OpenStudy (nirmalnema):

@raechelvictoria do u get it

OpenStudy (anonymous):

@hartnn (x-(1-3i))

OpenStudy (anonymous):

@nirmalnema Isn't the other factor supposed to be (x-(1+3i))?

OpenStudy (nirmalnema):

yup its correct...

hartnn (hartnn):

x-(1-3i) = x-1+3i is what i showed and other will be x-(1+3i) = x-1-3i ok?

OpenStudy (nirmalnema):

(x-2)(x-(1+3i))(x-(1-3i)) (x-2)(x^2-2x-9i^2) x^3-2x^2-9i^2x-2x^2+4x-18i^2 now i^2=(-1) so you will get x^3-4x^2+13x+18=0

OpenStudy (nirmalnema):

i put the wrong sign...sorry

hartnn (hartnn):

there is an error on your second step, you left out the '+1'

hartnn (hartnn):

and @raechelvictoria , which step are you on ? stuck or trying ?

OpenStudy (anonymous):

@hartnn I'm on the step you just showed :)

hartnn (hartnn):

ok, so we have 3 factors as x-2, x-1+3i, x-1-3i so, your f(x) = (x-2)(x-1+3i)(x-1-3i) = .... ? can you simplify?

OpenStudy (anonymous):

@hartnn the part where i'm having trouble is the part where you simplify the complex numbers :(

hartnn (hartnn):

ok, we use the fact that (a+b)(a-b) = a^2-b^2 so, we take, (x-1+3i)(x-1-3i) = ... take a= x-1 and b = 3i \(\large [(x-1)+(3i)][(x-1)-(3i)]=.....? \\knowing, (a-b)(a+b)=a^2-b^2\)

OpenStudy (anonymous):

\[(x-1)^{2} + 9\] ?

hartnn (hartnn):

correct! good :) f(x) = (x-2)(x^2-2x+10) right ? simplify this further, this part is easy :)

OpenStudy (anonymous):

\[x ^{3}-4x ^{2}+14x-20?\]

hartnn (hartnn):

yes! absolutely correct :)

OpenStudy (anonymous):

@hartnn thank you so much! :)

hartnn (hartnn):

you're welcome ^_^